# Diophantine Graphs

### v1

@@ Abstract
This manuscript introduces Diophantine labeling, a new way of labeling of the vertices for finite simple undirected graphs with some divisibility condition on the edges. Maximal graphs admitting Diophantine labeling are investigated and their number of edges are computed. Some number-theoretic techniques are used to characterize vertices of maximum degree and nonadjacent vertices. Some necessary and sufficient conditions for vertices of equal degrees are found.
Keywords: Graph labeling, Prime graph, Diophantine graph, $p$-adic valuation.
MSC code: 05A10, 05C07, 05C78, 11A05, 11A25, 11B75, 11D04, 11D88.
@@ 1 Introduction
A graph labeling is an assignment of real values to the vertices or edges that meet some conditions. The definition of prime labeling was originated by R. Entringer and introduced formally in a paper by R. Tout [12]. A graph $G=(V,E)$ with $n$ vertices has a prime labeling if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for every $uv\in E$, $(f(u),f(v))=1$. Many results of prime labeling and other types of prime labeling can be found in the dynamic survey of graph labeling by J.A. Gallian [6]. Other variants of prime labeling have been introduced by many researchers (see for example [4], [5], [8], [13]).
This paper gives a generalization for the concept of prime graphs. More precisely, we define a graph $G$ with $n$ vertices to be a Diophantine graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for every $uv\in E$, $(f(u),f(v))\mid n$. Clearly, any prime graph is a Diophantine graph. Moreover, any Diophantine graph with a prime number of vertices is a prime graph. In the maximal Diophantine graph with $n$ vertices, an explicit formulas of the number of edges and the degree of a vertex are proved, the full degree vertices are determined. Furthermore, we provide sufficient and necessary conditions that determine the equality of degrees among non-full degree vertices in the maximal Diophantine graphs.
The paper is organized as follows. Section 2 includes some definitions, preliminaries, lemmas and direct observations. Section 3 presents the number of edges formula in the maximal Diophantine graph with $n$ vertices. Section 4 is partitioned into three subsections, each unveiling important results related to the maximal Diophantine graph with order $n$. Subsection 4.1 covers a characterization of labels of vertices with full degree and a formula that compute the degrees of any vertex. Subsection 4.2 and 4.3 encompass some necessary and sufficient conditions for equality of degrees.
@@ 2 Definitions and Preliminaries
In the beginning of this section, we give the definitions of the prime graphs and Diophantine graphs. It also explores the relation between them and some direct observations.
@@
[11], [12] Let $G$ be a graph with $n$ vertices. The graph $G$ is called a prime graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for all $uv\in E$, $(f(u),f(v))=1$. Such a map $f$ is called a prime labeling of $G$. The maximal prime graph with $n$ vertices, denoted by $R_{n}$, is a prime graph such that adding any new edge yields a non-prime graph.
@@
Let $G$ be a graph with $n$ vertices. The graph $G$ is called a Diophantine graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for all $uv\in E$, $(f(u),f(v))\mid n$. Such a map $f$ is called a Diophantine labeling of $G$. The maximal Diophantine graph with $n$ vertices, denoted by $D_{n}$, is a Diophantine graph such that adding any new edge yields a non-Diophantine graph.
@@
The Petersen graph $P$ is a Diophantine graph, but the complete graph $K_{5}$ is not.
66662222333344441010101011117777888899995555
11112222333344445555
@@
It is easy to see that every prime graph is a Diophantine graph, but the converse is not necessarily true, e.g., $D_{4},D_{6},D_{8}$. However, in the case that $n$ is a prime number, a graph is Diophantine if and only if it is a prime graph. Moreover, $D_{n}\cong R_{n}.$ The proof is straightforward using the following lemma.
@@
Let $n>1$. Then we have $n$ is a prime number  if and only if  for all $a,b\leq n$, $a\neq b$, [$(a,b)=1$ if and only if $(a,b)\mid n$].
Proof.
Let $n$ be a prime number, $a,b<n$ and $a\neq b$. If $(a,b)=1,$ then $(a,b)\mid n$. On the other hand, if $(a,b)\mid n,$ then either $(a,b)=1$ or $(a,b)=n$. In the case that $(a,b)=1$, we have nothing to proof. In the other case that $(a,b)=n,$ we get $a=b=n$ which is a contradiction with the hypothesis.
Conversely, Assume $a\neq b$ and [$(a,b)=1$ if and only if $(a,b)\mid n$] for all $a,b<n$. Suppose, by the way of contradiction, $n>1$ is a composite number. Then there exists prime $p\mid n$ such that $p\leq\frac{n}{2}$. Let $a=p\leq n$, $b=2p\leq n$. Hence, $(a,b)=p\mid n$ and $(a,b)\neq 1,$ which contradicts the assumption. Hence, the result follows. ∎
In the subsequent sections, in any graph $G$ of order $n$, we call $f^{*}(u):=\frac{f(u)}{(f(u),n)}$ the reduced label of a vertex $u$ and we denote the set of all multiples of a positive integer $a$ up to $n$ by $M_{a}$, i.e. $M_{a}:=\{ta\ :\ t=1,2,\dots,\lfloor\frac{n}{a}\rfloor\}.$ Now, we present a crucial definition and an important number-theoretic lemmas that will allow us to formulate and prove our results smoothly.
Definition 2.3.
For a given positive integer $n$ and prime number $p$, the successor of the $p$-adic valuation is defined by $\acute{v}_{p}(n):=v_{p}(n)+1,$ where $v_{p}(n)$ is the $p$-adic valuation of an integer $n.$ The number $p^{\acute{v}_{p}(n)}$ is called the critical prime power number with respect to $p,n$.
Lemma 2.2.
Let $u,v\in V(D_{n})$ and $f$ be a Diophantine labeling of $D_{n}$. Then we have
  • (i)
@@
The proof of part $(i)$ is straightforward. Let us prove the second part. The $p$-adic valuation function [1] is extended normally to rational numbers as follows
@@
for any $a,b\in\mathbb{Z}^{+}$ and for every prime number $p.$ Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v).$ Using the basic properties of the $p$-adic valuation function, we have
@@
$v_{p}(n)<v_{p}(f(u)),\quad v_{p}(f(u))\leq v_{p}(n)\leq v_{p}(f(v))\quad\mbox{and}\quad v_{p}(f(v))<v_{p}(n).$
Hence,
$\frac{f(u)}{(f(u),n)}\ \Bigg{|}\ \frac{f(v)}{(f(v),n)},$
which completes the proof. ∎
In the proof of Lemma 2.3, we will use the following identity: $\frac{1}{j}-\frac{1}{j+1}=\left(1+\frac{2}{j}\right)\left[\frac{1}{j+1}-\frac{1}{j+2}\right]$, $j\in\mathbb{Z}^{+}$.
@@
$\begin{cases}\lfloor\sqrt{n}\rfloor\in\left(\frac{n}{\lfloor\sqrt{n}\rfloor+1},\frac{n}{\lfloor\sqrt{n}\rfloor}\right]\quad\mbox{and}\quad\frac{n}{\lfloor\sqrt{n}\rfloor}-\frac{n}{\lfloor\sqrt{n}\rfloor+1}<1,&\quad\mbox{if}\quad n<\lfloor\sqrt{n}\rfloor(\lfloor\sqrt{n}\rfloor+1);\\ &\\ \lfloor\sqrt{n}\rfloor\notin\left(\frac{n}{\lfloor\sqrt{n}\rfloor+1},\frac{n}{\lfloor\sqrt{n}\rfloor}\right]\quad\mbox{and}\quad\frac{n}{\lfloor\sqrt{n}\rfloor}-\frac{n}{\lfloor\sqrt{n}\rfloor+1}\geq 1,&\quad\mbox{if}\quad n\geq\lfloor\sqrt{n}\rfloor(\lfloor\sqrt{n}\rfloor+1).\\ \end{cases}$
@@
Let $a,b$ and $n\in\mathbb{N}$ such that $\lfloor\frac{n}{a}\rfloor=\lfloor\frac{n}{b}\rfloor$. Therefore, there exists $j\in\{1,2,\dots,n\}$ such that $\frac{n}{j+1}<a,b\leq\frac{n}{j}$. Consequently, $\frac{n}{j}-\frac{n}{j+1}>1$, since $a\neq b$. Using Lemma 2.3, we get that there exists $i\in\left\{\lfloor\sqrt{n}\rfloor,\lfloor\sqrt{n}\rfloor+1,\dots,n\right\}$ such that
@@
Since $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}},$ therefore Lemmas 2.2 part $(i)$ implies that
@@
Hence, the result follows. ∎
@@
The proof of Lemma 2.6 is a direct application of the fact $\left\lfloor\frac{x}{n}\right\rfloor=\left\lfloor\frac{\lfloor x\rfloor}{n}\right\rfloor$ for any $x\in\mathbb{R}$ and $n\in\mathbb{Z}^{+}.$ Lemmas 2.5 and 2.6 will be used in the proof of Theorem 4.5.
@@ 3 Number of Edges of $D_{n}$
The following lemma is a characterization of nonadjacent vertices in the maximal Diophantine graphs $D_{n}$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. For each $u,v\in V(D_{n}),$ $uv\notin E(D_{n})$ if and only if there exists a prime number $p$ such that
$f(u),f(v)\in M_{p^{\acute{v}_{p}(n)}}:=\left\{kp^{\acute{v}_{p}(n)}:\ k=1,2,\dots,\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\right\}.$
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and let $u,v\in V(D_{n})$. Then we have
$\begin{split}uv\notin E(D_{n})&\quad\mbox{if and only if}\quad(f(u),f(v))\nmid n\\ &\quad\mbox{if and only if}\quad v_{p}\left(\big{(}f(u),f(v)\big{)}\right)>v_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad\min\left\{v_{p}(f(u)),\ v_{p}(f(v))\right\}\geq v_{p}(n)+1=\acute{v}_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad v_{p}(f(u))\geq\acute{v}_{p}(n)\quad\mbox{and}\quad v_{p}(f(v))\geq\acute{v}_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad p^{\acute{v}_{p}(n)}\mid f(u)\quad\mbox{and}\quad p^{\acute{v}_{p}(n)}\mid f(v)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad f(u),\ f(v)\in M_{p^{\acute{v}_{p}(n)}}\quad\mbox{for some prime number $p$.}\hskip 139.41832pt\square\end{split}$
@@
The number of edges of the maximal Diophantine graph $D_{n}$ with order $n$ is given by
$C(n,2)-\sum_{1\leq i\leq\pi(n)}C\Bigg{(}\Bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg{\rfloor},2\Bigg{)}+\sum_{1\leq i<j\leq\pi(n)}C\Bigg{(}\Bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}}(n)p_{j}^{\acute{v}_{p_{j}}(n)}}\Bigg{\rfloor},2\Bigg{)}-\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right),$
where $p_{1},p_{2},\dots,p_{\pi(n)}<n$ are distinct prime numbers, $\pi(n)$ is the number of prime numbers not exceeding $n$ and $C(n,r)$ is the binomial coefficient with $C(n,r):=0$ if $n<r$ and $C(n,r):=\binom{n}{r}$ if $n\geq r$.
Proof. Let $D_{n}$ be the maximal Diophantine graph with $n$ vertices, $f$ be a Diophantine labeling of $D_{n}$ and $p_{i}<n,\ i=1,2,\dots,\pi(n)$ be distinct prime numbers. Using Lemma 3.1, we get for each $u,v\in V(D_{n})$, $uv\notin E(D_{n})$ if and only if $f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}}$ for some $i=1,2,\dots,\pi(n)$. Define
$S_{p_{i}}:=\big{\{}\{u,v\}\subset V(D_{n}):f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}},\ f(u)\neq f(v)\big{\}},$
$i=1,2,\dots,\pi(n)$. So, we have the following equation
@@
$|E(D_{n})|=C(n,2)-\left[\sum_{1\leq i\leq\pi(n)}|S_{p_{i}}|-\sum_{1\leq i<j\leq\pi(n)}|S_{p_{i}}\cap S_{p_{j}}|+\dots+(-1)^{\pi(n)-1}\left|\bigcap\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|\right].$
Therefore,
$\begin{split}|E(D_{n})|&=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\cap M_{p_{j}^{\acute{v}_{p_{j}}(n)}}\right|,2\right)-\\ &\qquad\qquad\quad\dots+(-1)^{\pi(n)}C\left(\left|\bigcap\limits_{1\leq i\leq\pi(n)}M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right).\\ &=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\right\rfloor,2\right)-\\ &\qquad\qquad\quad\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right).\end{split}$

### v2

# Linear Diophantine Graphs

@@ Abstract
This manuscript introduces linear Diophantine labelling, a new method for assigning labels to the vertices of finite, simple, undirected graphs. A key feature of this method is a divisibility condition imposed on the edges, incorporating number-theoretic properties into graph labelling. The study focuses on identifying maximal graphs that admit such labellings and computes their number of edges and the degree of each vertex. Number-theoretic techniques are employed to examine structural properties, including the characterization of maximum degree vertices and conditions for nonadjacent vertices. The manuscript also establishes necessary and sufficient conditions for vertices with equal degrees, offering new insights into the interaction between graph theory and number theory.
Keywords: Graph labelling, Prime graph, Diophantine graph, Equal degrees, $p$-adic valuation.
AMS Subject Classification: 05C07, 05C78, 11A05, 11A25, 11D88.
@@ 1 Introduction
A graph labelling is an assignment of real values to the vertices or edges or both edges and vertices that meet some conditions. It constitutes a significant area within graph theory and has numerous applications, including in coding theory, radar systems, astronomy, missile guidance, communication networks, cryptography, network security, X-ray crystallography, and database management. In this manuscript, we investigate a new type of labeling, referred to as linear Diophantine labeling, which constitutes a generalization of prime labeling. Labeling problems are pivotal in studying computational complexity and algorithmic graph theory, serving as a bridge between number theory and graph theory. The concept of prime labelling was first introduced by R. Entringer and later formally defined in a paper by R. Tout [12]. A simple graph $G=(V,E)$ with $n$ vertices is said to have a prime labelling if there is a bijective function $f$ mapping $V$ to $\{1,2,\dots,n\}$ in which for every $uv\in E$, the values $f(u),f(v)$ are relatively prime, where $V$ is the vertex set and $E$ is the edge set. Many results of prime labelling and other types of prime labelling can be found in the dynamic survey of graph labelling by J.A. Gallian [6]. Other variants of prime labelling have been introduced by many researchers (see, for example, [4], [5], [8], [14]).
This paper gives a generalization for the concept of prime graphs. More precisely, a simple graph $G$ with $n$ vertices is defined as a linear Diophantine graph (or simply, a Diophantine graph) if there is a bijective function $f$ mapping $V$ to $\{1,2,\dots,n\}$ in which for every $uv\in E$, the greatest common divisor of $f(u)$ and $f(v)$ divides $n$. Clearly, any prime graph is a Diophantine graph. Moreover, any Diophantine graph with a prime number of vertices is a prime graph. In a maximal Diophantine graph with $n$ vertices, an explicit formula of the number of edges and the degree of a vertex are proved, the full degree vertices are determined. Furthermore, we provide sufficient and necessary conditions that determine the equality of degrees among vertices whose degrees are smaller than n1n-1 in a maximal Diophantine graphs.
The paper is organized as follows. Section 2 includes some definitions, preliminaries, lemmas and direct observations. Section 3 presents the number of edges formula in a maximal Diophantine graph with $n$ vertices. Section 4 is split into three subsections, each unveiling important results related to a maximal Diophantine graph with order $n$. Subsection 4.1 covers a characterization of labels of vertices with full degree and a formula that computes the degrees of any vertex. Subsection 4.2 and 4.3 encompass some necessary and sufficient conditions for equality of degrees.
@@
Theorem 1.1.
[3], [9] (Fundamental Theorem of Arithmetic)
Every positive integer $n>1$ can be expressed as a product of primes; this representation is unique, apart from the order in which the factors occur.
Definition 1.1.
[2] Two graphs $G_{1},G_{2}$ are said to be isomorphic, denoted by $G_{1}\cong G_{2}$, if there exists a bijective map $\varphi:V(G_{1})\rightarrow V(G_{2})$ such that for all $u,v\in V(G_{1})$,
$uv\in E(G_{1})\quad\mbox{if and only if}\quad\varphi(u)\varphi(v)\in E(G_{2}).$
@@ 2 Definitions and Preliminaries
In the beginning of this section, we present the definitions of the prime graphs and the Diophantine graphs and we explore the relation between them. Additionally, we provide direct observations on several number-theoretic lemmas.
@@
[11], [12] A simple graph $G$ of order $n$ is called a prime graph if there exists a bijective mapping $f:V\rightarrow\{1,2,\dots,n\}$ such that $(f(u),f(v))=1$ for all $uv\in E$. The function $f$ is called a prime labelling of $G$. A prime graph of order $n$ is called a maximal prime graph, denoted by $(R_{n},f)$, if any new edge added to this graph turns it into a non-prime graph.
@@
A simple graph $G$ of order $n$ is called a linear Diophantine graph (or simply, a Diophantine graph) if there exists a bijective mapping $f:V\rightarrow\{1,2,\dots,n\}$ such that $(f(u),f(v))\mid n$ for all $uv\in E$. The function $f$ is called a Diophantine labelling of $G$. A Diophantine graph of order $n$ is called a maximal Diophantine graph, denoted by $(D_{n},f)$, if any new edge added to this graph turns it into a non-Diophantine graph. If there is no ambiguity, we omit $f$ from $(D_{n},f)$ and simply write $D_{n}$.
@@
The Peterson graph $P$ is a Diophantine graph, but the complete graph $K_{5}$ is not.
6622334410101177889955
1122334455
@@
It is easy to see that every prime graph is a Diophantine graph, but the converse is not necessarily true, e.g., $D_{4},D_{6},D_{8}$.
Theorem 2.1.
Let $n>1$ be a prime number. A graph $G$ of order $n$ is a Diophantine graph if and only if $G$ is a prime graph.
Proof.
The proof is clear. ∎
Theorem 2.2.
For every $n>1$, $D_{n}\cong R_{n}$ if and only if $n$ is a prime number.
Proof.
Let $K_{n}$ be the complete graph of order $n$ and $f:V(K_{n})\rightarrow\{1,2,\dots,n\}$ be a bijective map. Define two sets
$E(R_{n}):=\{uv\in E(K_{n}):(f(u),f(v))=1\}$
and
$E(D_{n}):=\{uv\in E(K_{n}):(f(u),f(v))\mid n\}.$
Let $D_{n}\cong R_{n}$. Then $E(R_{n})=E(D_{n})$. Therefore,
$(f(u),f(v))\mid n\quad\mbox{if and only if}\quad(f(u),f(v))=1.$
Suppose, by the way of contradiction, $n>1$ is a composite number. Then there exists a prime number $p\mid n$ such that $p\leq\frac{n}{2}$. Let $f(u)=p\leq n$ and $f(v)=2p\leq n$. Then $(f(u),f(v))=p\mid n$ and $(f(u),f(v))\neq 1$ which contradicts $(f(u),f(v))=1$. Hence, $n$ is a prime number.
Conversely, let $n$ be a prime number and let $uv\in E(R_{n})$. Then $(f(u),f(v))=1$. Consequently, $(f(u),f(v))\mid n$. Therefore, $uv\in E(D_{n})$. On the other hand, let $uv\in E(D_{n})$. Then $(f(u),f(v))\mid n$. Since $n$ is a prime number, we have either $(f(u),f(v))=1$ or $(f(u),f(v))=n$. In the case that $(f(u),f(v))=1$, we have nothing to proof. In the other case that $(f(u),f(v))=n,$ we get $f(u)=f(v)=n$ which is a contradiction with the assumption. Consequently, $(f(u),f(v))=1$. Therefore, $uv\in E(R_{n})$. Hence, $D_{n}\cong R_{n}$. ∎
In the subsequent sections, the $p$-adic valuation of $n\in\mathbb{Z}^{+}$ is defined by $v_{p}(n):=t$ such that $p^{t}\mid n$ and $p^{t+1}\nmid n$, where $p$ is a prime number and $t\in\mathbb{N}:=\mathbb{Z}^{+}\cup\{0\}$ and $\mathbb{Z}^{+}$ is the set of positive integers. The set of all multiple of $a\in\mathbb{Z}^{+}$ up to $n$ is denoted by $M_{a}$, i.e. $M_{a}:=\left\{ka\ :\ k=1,2,\dots,\left\lfloor\frac{n}{a}\right\rfloor\right\}.$ In any graph $G$ of order $n$, we call $f^{*}(u):=\frac{f(u)}{(f(u),n)}$ the reduced label of a vertex $u$ and the neighborhood of $s\in V(G)$ is defined by $N(s):=\{u\in V(G):su\in E(G)\}$. Now, we present a crucial definition and important number-theoretic lemmas that will allow us to formulate and prove our results smoothly.
Definition 2.3.
A number $p^{\acute{v}_{p}(n)}$ is called a critical prime power number with respect to a prime number $p$ and a positive integer $n$, where $\acute{v}_{p}(n):=v_{p}(n)+1$ is called the successor of the $p$-adic valuation.
@@
Let $u,v$ be two vertices in $D_{n}$ and $f$ be a Diophantine labelling of $D_{n}$. Then
  1. i.
@@
The proof of part $i.$ is straightforward. Let us prove the second part. The $p$-adic valuation function [1] is extended normally to rational numbers as follows
@@
for any $a,b\in\mathbb{Z}^{+}$ and for every prime number $p$. Let $u,v$ be two vertices in $V(D_{n})$ such that $f(u)\mid f(v).$ Using the basic properties of the $p$-adic valuation function, we have
@@
$v_{p}(n)<v_{p}(f(u)),\ v_{p}(f(u))\leq v_{p}(n)\leq v_{p}(f(v))\ \mbox{and}\ v_{p}(f(v))<v_{p}(n).$
Therefore,
$\frac{f(u)}{(f(u),n)}\ \Bigg|\ \frac{f(v)}{(f(v),n)}.$
Hence, the proof follows. ∎
We will use the following identity: $\frac{1}{j}-\frac{1}{j+1}=\left(1+\frac{2}{j}\right)\left[\frac{1}{j+1}-\frac{1}{j+2}\right]$, $j\in\mathbb{Z}^{+}$ for the proof of Lemma 2.2.
@@
[…]
@@
Clearly, the statement is true in case of positive integers $a$ or $b$ are greater than or equal to $n\in\mathbb{Z}^{+}$. Now, let $a,b<n$ such that $\lfloor\frac{n}{a}\rfloor=\lfloor\frac{n}{b}\rfloor$. Therefore, there exists $j\in\{1,2,\dots,n\}$ such that $\frac{n}{j+1}<a,b\leq\frac{n}{j}$. Consequently, $\frac{n}{j}-\frac{n}{j+1}>1$, since $a\neq b$. Using Lemma 2.2, we get that there exists $i\in\left\{\lfloor\sqrt{n}\rfloor,\lfloor\sqrt{n}\rfloor+1,\dots,n\right\}$ such that
@@
Since $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}},$ therefore Lemmas 2.1, part $i.$, implies that
@@
Hence, the proof follows. ∎
@@
The proof of Lemma 2.5 is a direct application of the fact $\left\lfloor\frac{x}{n}\right\rfloor=\left\lfloor\frac{\lfloor x\rfloor}{n}\right\rfloor$ for any $x\in\mathbb{R}$ and $n\in\mathbb{Z}^{+},$ where $\mathbb{R}$ is the set of real numbers. Lemmas 2.4 and 2.5 will be used in the proof of Theorem 4.6.
@@ 3 Number of Edges of $D_{n}$
In this section, we derive a formula for the number of edges of maximal Diophantine graphs $D_{n}$ using a lemma that characterizes nonadjacent vertices. In the remainder of this paper, let $D_{n}$ denote a maximal Diophantine graph with order $n$ that admits a Diophantine labelling $f$.
@@
For each two vertices $u,v$ in $D_{n}$, $uv\notin E(D_{n})$ if and only if there exists a prime number $p$ such that $f(u),f(v)\in M_{p^{\acute{v}_{p}(n)}},$ where $M_{p^{\acute{v}_{p}(n)}}$ is the set of multiple of $p^{\acute{v}_{p}(n)}$ up to $n$.
Proof. Let $u,v$ be two vertices in $D_{n}$ such that
[…]
@@
The size of $D_{n}$ is given by
$\begin{split}|E(D_{n})|=C(n,2)&-\sum_{1\leq i\leq\pi(n)}C\Bigg(\Bigg\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg\rfloor,2\Bigg)+\sum_{1\leq i<j\leq\pi(n)}C\Bigg(\Bigg\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}}(n)p_{j}^{\acute{v}_{p_{j}}(n)}}\Bigg\rfloor,2\Bigg)\\ &-\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right),\end{split}$
where $p_{i}<n$, $i=1,2,\dots\pi(n)$ are distinct prime numbers, $\pi(n)$ is the number of prime numbers not exceeding $n$ and $C(n,r)$ is the binomial coefficient, defined as $C(n,r):=0$ if $n<r$ and $C(n,r):=\binom{n}{r}$ if $n\geq r$.
Proof. Let $p_{i}<n$, $i=1,2,\dots,\pi(n)$ be distinct prime numbers. Then, using Lemma 3.1, we get for each two vertices $u,v$ in $D_{n}$, $uv\notin E(D_{n})$ if and only if $f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}}$ for some $i=1,2,\dots,\pi(n)$. Define
$S_{p_{i}}:=\left\{uv\notin E(D_{n}):f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\},\quad i=1,2,\dots,\pi(n).$
Thus, we have the following equation
@@
$\begin{split}|E(D_{n})|&=C(n,2)-\sum_{1\leq i\leq\pi(n)}|S_{p_{i}}|+\sum_{1\leq i<j\leq\pi(n)}|S_{p_{i}}\cap S_{p_{j}}|-\dots+(-1)^{\pi(n)}\left|\bigcap\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|.\\ &=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\cap M_{p_{j}^{\acute{v}_{p_{j}}(n)}}\right|,2\right)\\ &\qquad\qquad\quad-\dots+(-1)^{\pi(n)}C\left(\left|\bigcap\limits_{1\leq i\leq\pi(n)}M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right).\\ &=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right)+\hskip-2.84544pt\sum_{1\leq i<j\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\right\rfloor,2\right)\\ &\qquad\qquad\quad-\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right).\hskip 116.65646pt\square\end{split}$